. Out of a 940 Km journey, a part of the journey was covered by a motor car at a speed of
72 Km/hr. The remaining part of the journey was covered by train at the speed of 84
Km/hr. If the total distance was covered in 12 hours, find the distance traveled by the
train and the time taken by it to cover that distance.

Answer by stanbon(75887)

Out of a 940 Km journey, a part of the journey was covered by a motor car at a speed of 72 Km/hr.
Car DATA::
rate = 72 km/hr ; distance = x km ; time = d/r = x/72 hrs
-------------------------------------
The remaining part of the journey was covered by train at the speed of 84
Km/hr.
Train DATA:
rate = 84 km/hr ; distance = 940-x km ; time = d/r = (940-x)/84 hrs
---------------------------
If the total distance was covered in 12 hours, find the distance traveled by the train and the time taken by it to cover that distance.
Equation:
time + time = 12 hrs
x/72 + (940-x)/84 = 12
etc.
x = 408 km
940-x = 532 km (train distance)
532/84 = 6 1/3 hrs (train time)
-----------------------------------
Cheers,
Stan H.
-------------------
----------